# Task-agnostic Pre-training and Task-guided Fine-tuning for Versatile Diffusion Planner

### v1

@@ 4 Related Work
Diffusion Models in RL. Diffusion models are a leading class of generative models, achieving state-of-the-art performance across a variety of tasks, such as image generation (Ramesh et al., 2021), audio synthesis (Kong et al., 2020; Huang et al., 2023), and drug design (Schneuing et al., 2022; Guan et al., 2024). Recent studies have applied them in imitation learning to model human demonstrations and predict future actions (Li et al., 2024; Reuss et al., 2023). Other approaches have trained conditional diffusion models either as planners (Ajay et al., 2022; Brehmer et al., 2024) or policies (Hansen-Estruch et al., 2023; Kang et al., 2024). However, most of these efforts focus on single-task settings. While some recent works aim to extend diffusion models to multi-task scenarios, they often rely on additional conditions, such as prompts (He et al., 2024) or preference labels (Yu et al., 2024). These methods are limited by their dependence on expert data or explicit task knowledge. In contrast, our method learns broad action-sequence distributions from inferior data to enhance action priors, enabling effective generalization across a range of downstream tasks.
Fine-tuning Diffusion Models. Despite the impressive success of diffusion models, they often face challenges in aligning with specific downstream objectives, such as image aesthetics (Schuhmann et al., 2022), fairness (Shen et al., 2023), or human preference (Xu et al., 2024), primarily due to their training on unsupervised data. Some methods have been proposed to address this issue by directly fine-tuning models using downstream objectives (Prabhudesai et al., 2023; Clark et al., 2023), but they rely on differentiable reward models, which are impractical in RL since accurately modeling rewards with neural networks is quite costly (Kim et al., 2023). Other methods reformulate the denoising process as an MDP and apply policy gradients for fine-tuning (Black et al., 2023; Fan et al., 2024). However, they heavily depend on strong pre-trained models and have proven ineffective in our case. Our goal is to fine-tune a less powerful model that has been trained on inferior data.
Concurrent with our work, DPPO (Ren et al., 2024) also explores reward fine-tuning for refining RL diffusion planners. However, their approach focuses exclusively on single-task settings and allows access to expert demonstrations. In contrast, we train our model on multi-task data without the need for superior demonstrations. Additionally, we analyze the limitations of current regularization methods for versatile RL diffusion models and propose a new regularizer that improves the performance of sub-optimal pre-trained models.
@@ 5 Experiments
In this section, we conduct experiments to evaluate our proposed method and address the following questions: (1) How does SODP’s performance compare to current methods? (2) Can SODP scale to high-dimensional observation inputs? (3) How does SODP achieve higher rewards during online fine-tuning?
@@ 5.1 Experimental Setup
We evaluate SODP in both state-based and image-based environments. For state-based environments, we conduct experiments using the Meta-World benchmark (Yu et al., 2019). For image-based environments, we perform experiments using the Adroit benchmark (Rajeswaran et al., 2017).
@@
Meta-World. The Meta-World benchmark comprises 50 distinct manipulation tasks, each requiring a Sawyer robot to interact with various objects. These tasks are designed to assess the robot’s ability to handle different scenarios, such as grasping, pushing, pulling, and manipulating objects of varying shapes, sizes, and complexities. While the state space and reward functions differ across tasks, the action space remains consistent. Following recent studies (He et al., 2024; Hu et al., 2024), we extend all tasks to a random-goal setting, referred to as MT50-rand.
Adroit. The Adroit benchmark includes three dexterous manipulation tasks, requiring a 24-degree-of-freedom dexterous hand to solve complex challenges such as in-hand manipulation and tool use. The goals in this environment are also randomized. For Adroit, we use images as the observation to assess whether our method can scale to high-dimensional input.
@@
Datasets. Following previous work (He et al., 2024), for Meta-World, we use a sub-optimal dataset comprising the first 50% of experiences (50M transitions) obtained from the replay buffer of a SAC (Haarnoja et al., 2018) agent during training. To verify the applicability of our method to tasks of varying difficulty levels, we divide the entire dataset into four subsets based on the task categories presented in Seo et al. (2023). All baselines and our pre-training stage are trained on sub-optimal datasets. We then use 1M online transitions per task for fine-tuning. For Adroit, we train a VRL3 (Wang et al., 2022a) agent for each task and use the initial 30% experiences (90K transitions) from the converged replay buffer. All baselines are trained on expert demonstrations and our pre-training stage is trained on sub-optimal transitions.
@@
Baselines. For Meta-World, we compare our proposed SODP with the following baselines: (1) MTBC.Extended BC to multi-task offline policy learning through network scaling and a task-ID-conditioned actor. (2) MTIQL.Extended IQL (Kostrikov et al., 2021) with multi-head critic networks and a task-ID-conditioned actor for multi-task policy learning. (3) MTDT.Extended Decision Transformer (DT) (Chen et al., 2021a) to multitask settings by incorporating task ID encoding and state inputs for task-specific learning. (4) Prompt-DT(Xu et al., 2022). An extension of DT, which is designed to leverage multi-task data to generalize policies for previously unseen tasks. It generates actions by utilizing trajectory prompts and reward-to-go signals. (5) MTDIFF(He et al., 2024). A diffusion-based approach that integrates Transformer architectures with prompt learning to facilitate generative planning in multitask offline environments. (6) HarmoDT(Hu et al., 2024). A DT-based approach that leverages parameter sharing to exploit task similarities while mitigating the adverse effects of conflicting gradients simultaneously. The results for these baselines are directly replicated from those reported in HarmoDT (Hu et al., 2024).
Learning policies for Adroit is more challenging due to the significantly larger state and action spaces. The baselines used in Meta-World struggled to handle this high-dimensional data structure. Therefore, we compare SODP with following baselines designed for complex environments: (1) BCRNN(Mandlekar et al., 2021). A variant of BC that employs a Recurrent Neural Network (RNN) as the policy network, predicting the sequence of actions based on the sequence of states as input. (2) IBC(Florence et al., 2022). Extended BC with energy-based models (EBM) to train implicit behavioral cloning policies. (3) Diffusion Policy(Chi et al., 2023). A diffusion-based approach that predicts future action sequences based on historical states. (4) DP3(Ze et al., 2024). A visual imitation learning algorithm that incorporates 3D visual representations into diffusion policies, using a point clouds encoder to process visual observations into visual features. The results for these baselines are directly replicated from those reported in DP3 (Ze et al., 2024).
5.2 Results
@@
We use the average success rate across all tasks as the evaluation metric and report the mean and standard deviation of success rates across three seeds. All baselines are trained on sub-optimal data. As shown in Table 1, our method achieves over a 60% success rate when learning from inferior data, outperforming all baseline methods. Compared to the existing state-of-the-art approach,
@@
our method demonstrates a 5.9% improvement. Notably, when compared to MTDIFF, the current leading method based on diffusion models, our approach shows a 24.4% improvement. MTBC performs the worst, as imitation learning heavily depends on data quality, and directly cloning behaviors from sub-optimal data typically results in inferior performance. In contrast, our method models versatile action distributions from low-quality data and leverages them as priors to guide policy optimization in downstream tasks, leading to improved performance.
To further analyze the learning dynamics, we sample 10 tasks and present their learning curves of SODP alongside two leading baselines, MTDIFF and HarmoDT, across five seeds. As shown in Figure 4, SODP rapidly converges to high success rates, surpassing the other two baselines. The pre-training stage equips the planner with comprehensive action distribution priors and allows it to rapidly transfer and enhance these capabilities across a variety of downstream tasks. As a result, the pre-training stage significantly accelerates convergence, leading to more efficient learning in the fine-tuning stage. The two baseline approaches struggle to address complex and challenging tasks such as basketball and hammer. In contrast, our method effectively guides the model to generate proficient actions, demonstrating the benefits of fine-tuning with policy gradient concerning return maximization. Moreover, while HarmoDT exhibits instability across different random seeds, our method demonstrates robustness against randomness.
@@
Figure 4: Learning efficiency. We sample 10 tasks and present the learning curves of SODP, MTDIFF, and HarmoDT across five seeds. X-axis represents gradient steps. We pre-train the planner for $5e^{5}$ steps, followed by fine-tuning with a smaller number of steps. SODP rapidly converges to high success rates, whereas MTDIFF and HarmoDT struggle with some challenging tasks.
Does SODP generalize to high-dimensional observations? We scale our method to image-based
Adroit
[…]
BCRNN $0_{\pm 0}$ $0_{\pm 0}$ $9_{\pm 3}$ 3.0
BCRNN+3D $8_{\pm 14}$ $0_{\pm 0}$ $8_{\pm 1}$ 5,3
IBC $0_{\pm 0}$ $0_{\pm 0}$ $9_{\pm 2}$ 3.0
IBC+3D $0_{\pm 0}$ $0_{\pm 0}$ $10_{\pm 1}$ 3.3
Diffusion Policy $48_{\pm 17}$ $50_{\pm 5}$ $25_{\pm 4}$ 31.7
Simple DP3 $100_{\pm 0}$ $58_{\pm 4}$ $46_{\pm 5}$ 68.0
DP3 $100_{\pm 0}$ $62_{\pm 4}$ $43_{\pm 6}$ 68.3
SODP (ours) $67_{\pm 6}$ $96_{\pm 1}$ $59_{\pm 4}$ 73.9
observations using the Adroit benchmark by employing a point-cloud encoder from DP3 (Ze et al., 2024) to process the 3D scene represented by point clouds. Specifically, we capture depth images directly from the environment and convert them into point clouds using Open3D (Zhou et al., 2018). These point clouds are then processed by the DP3 Encoder, which maps them into visual features. We then train our diffusion planner following the same procedure in Algorithm 1 except the input states are visual features. Following DP3 (Ze et al., 2024), We compute the average of the highest 5 evaluation success rates during training and report the mean and std across 3 seeds. As shown in Table 2, our method outperforms the baseline approaches and achieves an 8.2% improvement, demonstrating its capability to scale up to high-dimensional inputs like images and point-clouds.
@@ 5.3 Effectiveness of BC Regularization
To demonstrate the effectiveness of our BC regularization, we conduct an ablation study on fine-tuning same pre-trained model with our BC regularization and other variants. We consider following variants:
  • SODP w/o regularization. This variant is similar to DDPO (Black et al., 2023) and DPPO (Ren et al., 2024), which fine-tunes the model directly using Eq. (12) without any regularization.
  • SODP_kl. This variant is similar to DPOK (Fan et al., 2024), with the addition of a KL regularization term to constrain the divergence between the fine-tuned model and the pre-trained model.
  • SODP_pl. This variant is similar to DLPO (Chen et al., 2024), incorporating the original diffusion pre-training loss (PL) into the fine-tuning objective to prevent the model from deviation.
The details of these variants are presented in Appendix D. Figure 5 demonstrates the effectiveness of our regularization in achieving a higher success rate. We observe that directly fine-tuning the model without any regularization results in the worst performance, with a decline in success rate, as the model may degrade the capabilities acquired from pre-training due to the lack of constraints. However, adding KL and PL is insufficient, as they cause oscillations near the pre-trained model. This aligns with the original intent of these regularizers, which is to prevent excessive deviation. This is reasonable for methods like DPOK and DLPO, which utilize pre-trained models such as Stable Diffusion (Rombach et al., 2022) and WaveGrad2 (Chen et al., 2021b). These models already exhibit strong generative capabilities without fine-tuning, and the goal is to make slight adjustments to align them with more fine-grained attributes, such as aesthetic scores and human preferences.
In contrast, our model is pre-trained on sub-optimal data and lacks the ability to solve complex tasks. We expect it to develop new skills for completing these tasks through fine-tuning. However, directly applying KL regularization to the pre-trained model leads to conservative policies that heavily rely on the existing capability, thereby confining the model to a sub-optimal region. While PL regularization allows some slight exploration, it is uncontrolled and random. Consequently, we observe that the KL regularization almost remains unchanged and the PL regularization slightly increases the performance in basketball but decreases in other tasks.
@@
Figure 5: Learning efficiency for different regularization. X-axis represents environment steps. Performance declines significantly without any regularization. Both KL and PL regularization confine the model to sub-optimal regions. In contrast, our BC regularization effectively guides the model away from these sub-optimal areas, facilitating the attainment of optimal actions.
Visualization.
We hypothesize that the effectiveness of our BC regularization lies in two aspects: (i) it ensures that our model can reuse the skills it has acquired, thereby preventing a decline in performance; (ii) It guides our model to effectively explore optimal regions due to the utilization of optimal $\mu$ as the target policy. To demonstrate this, we visualize trajectories of using the actions generated by our planner using t-SNE (Van der Maaten & Hinton, 2008). As shown in Figure 6, the trajectory distribution after fine-tuning with KL regularization closely resembles the original pre-training distribution, indicating that the model is reusing learned actions and lacks exploration into new regions. The exploration in PL is unstructured as it may lead to worse regions (e.g., the upper-left region in basketball). In contrast, our method demonstrates superior exploration capabilities to discover new, high-reward regions based on acquired knowledge (e.g., the lower-left region in basketball and the bottom region in plate-slide). Meanwhile, the model can derive valuable insights from pre-trained knowledge by exploiting discovered high-reward actions (e.g. the central region in plate-slide) while discarding low-reward actions.
@@
Figure 6: Visualization of trajectories using generated actions for different regularization. KL and PL regularization results in conservative policies with distributions closely resembling the original. Our BC regularization retains pre-trained knowledge while effectively discovering new actions that can lead to high rewards.
@@ 5.4 Effectiveness of Pre-training
We investigate the impact of pre-training. We compare the performance of SODP with a version trained from scratch (SODP_scratch). For SODP_scratch, we use the same rollouts generated by the pre-trained model to approximate the target policy and initialize the replay buffer.
Figure 7 shows that fine-tuning the planner from scratch results in worse performance. Without pre-training, the planner lacks an action prior to guide its behavior, leading to stagnation as it struggles to move towards high reward regions. Additionally, it becomes unstable, as the limited useful knowledge is easily disrupted by a large number of ineffective trials.

### v3

@@ 4 Related Work
Diffusion Models in RL. Diffusion models are a leading class of generative models, achieving state-of-the-art performance across a variety of tasks, such as image generation (Ramesh et al., 2021), audio synthesis (Kong et al., 2021; Huang et al., 2023), and drug design (Schneuing et al., 2024; Guan et al., 2023). Recent studies have applied them in imitation learning to model human demonstrations and predict future actions (Li et al., 2024; Reuss et al., 2023; He et al., 2024). Other approaches have trained conditional diffusion models either as planners (Ajay et al., 2023; Brehmer et al., 2024; Yuan et al., 2024) or policies (Hansen-Estruch et al., 2023; Kang et al., 2024). However, most of these efforts focus on single-task settings. While some recent works aim to extend diffusion models to multi-task scenarios, they often rely on additional conditions, such as prompts (He et al., 2023) or preference labels (Yu et al., 2024). These methods are limited by their dependence on expert data or explicit task knowledge. In contrast, our method learns broad action-sequence distributions from inferior data to enhance action priors, enabling effective generalization across a range of downstream tasks.
Fine-tuning Diffusion Models. Despite the impressive success of diffusion models, they often face challenges in aligning with specific downstream objectives, such as image aesthetics (Schuhmann et al., 2022), fairness (Shen et al., 2024), or human preference (Xu et al., 2024), primarily due to their training on unsupervised data. Some methods have been proposed to address this issue by directly fine-tuning models using downstream objectives (Prabhudesai et al., 2023; Clark et al., 2024), but they rely on differentiable reward models, which are impractical in RL since accurately modeling rewards with neural networks is quite costly (Kim et al., 2023). Other methods reformulate the denoising process as an MDP and apply policy gradients for fine-tuning (Black et al., 2024; Fan et al., 2024). However, they heavily depend on strong pre-trained models and have proven ineffective in our case. Our goal is to fine-tune a less powerful model that has been trained on inferior data.
Concurrent with our work, DPPO (Ren et al., 2025) also explores reward fine-tuning for refining diffusion planners. However, their approach focuses on single-task settings and allows access to expert demonstrations. In contrast, we train our model on multi-task data without the need for superior demonstrations. Additionally, we analyze the limitations of current regularization methods for versatile RL diffusion models and propose a new regularizer that improves the performance of sub-optimal pre-trained models.
@@ 5 Experiments
In this section, we conduct experiments to evaluate our proposed method and address the following questions: (1) How does the performance of SODP compare to other methods that combine offline pre-training with online fine-tuning? (2) How does the performance of SODP compare to other multi-task learning approaches? (3) How does SODP achieve higher rewards during online fine-tuning?
@@ 5.1 Experimental Setup
We evaluate SODP in both state-based and image-based environments. We conduct experiments on the Meta-World benchmark (Yu et al., 2019) for both state-based and image-based tasks. We also perform image-based experiments on the Adroit benchmark (Rajeswaran et al., 2018). Image-based experimental results are presented in Appendix C.4.
@@
Meta-World. The Meta-World benchmark comprises 50 distinct manipulation tasks, each requiring a Sawyer robot to interact with various objects. These tasks are designed to assess the robot’s ability to handle different scenarios, such as grasping, pushing, pulling, and manipulating objects of varying shapes, sizes, and complexities. While the state space and reward functions differ across tasks, the action space remains consistent. Following recent studies (He et al., 2023; Hu et al., 2024), we extend all tasks to a random-goal setting, referred to as MT50-rand.
@@
Datasets. Following previous work (He et al., 2023), we use a sub-optimal offline dataset containing 1M transitions for each task. The dataset consists of the first 50% of experiences collected from the replay buffer of an SAC (Haarnoja et al., 2018) agent during training. To verify the applicability of our method to tasks of varying difficulty levels, we divide the entire dataset into four subsets based on the task categories presented in Seo et al. (2023).
Baselines. We compare our proposed SODP with several RL baselines that integrate offline pre-training with online fine-tuning: (1) Cal-QL(Nakamoto et al., 2024). Employs conservative offline initializations to facilitate online fine-tuning. (2) IBRL(Hu et al., 2023). Pre-trains a policy using imitation learning and subsequently fine-tunes it as a RL policy. (3) RLPD(Ball et al., 2023). Extends SAC by utilizing a sample buffer that incorporates both offline and online data. All baselines, along with SODP, are pre-trained on the same dataset containing 50M transitions and are subsequently fine-tuned on each task with 1M transitions.
@@
Besides, we compare SODP with the following multi-task RL baselines: (1) MTSAC.Extended SAC with one-hot task ID as additional input. (2) MTBC.Extended BC to multi-task learning through network scaling and a task-ID-conditioned actor. (3) MTIQL.Extended IQL (Kostrikov et al., 2022) with multi-head critic networks and a task-ID-conditioned actor for multi-task policy learning. (4) MTDQL.Extended Diffusion-QL (Wang et al., 2023) which is similar to MTIQL. (5) MTDT.Extended Decision Transformer (DT) (Chen et al., 2021a) to multitask settings by incorporating task ID encoding and state inputs for task-specific learning. (6) Prompt-DT(Xu et al., 2022). An extension of DT, which generates actions by utilizing trajectory prompts and reward-to-go signals. (7) MTDIFF(He et al., 2023). A diffusion-based approach that integrates Transformer architectures with prompt learning to facilitate generative planning in multitask offline environments. We extend it with a visual extractor in image-based Meta-World experiments. (8) HarmoDT(Hu et al., 2024). A DT-based approach that leverages parameter sharing to exploit task similarities while mitigating the adverse effects of conflicting gradients simultaneously. All baselines are trained on the 50M offline transitions.
5.2 Main Results
@@
We use the average success rate across all tasks as the evaluation metric and report the mean and standard deviation of success rates across three seeds. As shown in Table 1, our method achieves a 60.56% success rate when learning from inferior data, outperforming all baseline methods. Current offline-online RL methods face challenges in effectively learning from complex, low-quality offline data, resulting in insufficient priors to guide the online fine-tuning. Compared to the existing state-of-the-art multi-task approach, our method demonstrates a 5.9% improvement. Notably, when compared to MTDIFF, the current leading method based on diffusion models, our approach shows a 24.4% improvement.
@@
MTBC performs the worst, as imitation learning heavily depends on data quality, and directly cloning behaviors from sub-optimal data typically results in inferior performance. In contrast, our method models versatile action distributions from low-quality data and leverages them as priors to guide policy optimization in downstream tasks, leading to improved performance.
@@
Figure 4: Learning dynamics. We sample 10 tasks and present the learning curves of SODP, Cal-QL, IBRL and RLPD across five seeds. X-axis represents environment steps. SODP converges to higher success rates, whereas others struggle with solving the downstream task.
To further analyze the learning dynamics, we sample 10 tasks and present the learning curves of SODP alongside three offline-online baselines. As shown in Figure 4, SODP demonstrates a higher success rate after fine-tuning, whereas other baseline methods struggle to address the specific task. This performance disparity can be attributed to the inadequate extraction of meaningful guidance during the offline phase. Due to the presence of numerous low-reward trajectories in the sub-optimal dataset, pre-training on such data may misguide the value function. Furthermore, employing a Gaussian policy for behavior cloning fails to capture the multimodal action distributions embedded in the pre-training data. Consequently, this bias can affect the updates during online fine-tuning, potentially trapping the model in a sub-optimal plateau or even degrading its overall performance. The learning curves of SODP, in comparison to other multi-task baselines, are provided in Appendix C.6. We also conduct additional experiments on offline-online baselines, including both diffusion-based and transformer-based methods, with the results presented in Appendix C.3.
5.3 Efficiency Validation with Limited Online Samples
To validate the learning efficiency of SODP, we conduct experiments in an extreme scenario where the online fine-tuning steps are reduced to 100k per task. We compare SODP with three offline-online baselines, each using only 100k fine-tuning steps, as well as two leading multi-task RL baselines: MTDIFF and HarmoDT. Since MTDIFF and HarmoDT are purely offline algorithms, we collect online interaction samples during the SODP fine-tuning process. These samples are incorporated as a supplementary dataset alongside the original data, increasing the dataset size from 50M to 50M + 100k×$\times$50. We then train both MTDIFF and HarmoDT on this augmented dataset to ensure consistent data usage across our method and the baselines. As shown in Table 2, the performance of our method does not degrade significantly despite a dramatic reduction in fine-tuning steps, demonstrating the strong efficiency of SODP in learning from limited online transitions. In contrast, all three offline-online baselines experience a substantial decline due to insufficient online samples for correcting pre-trained knowledge. For multi-task RL baselines, HarmoDT remains stable when trained on the augmented dataset, whereas the performance of MTDIFF declines markedly. This may be attributed to the increased presence of inferior data collected during the online interaction phase. The generation process of MTDIFF is guided and subsequently biased by these newly introduced low-reward trajectories, whereas the relatively small proportion of low-reward online data compared to the original offline data does not affect the established parameter subspaces in HarmoDT.
Table 2: Average success rate of fine-tuning with 100k steps for offline-online methods and training on augmented sub-optimal data for multi-task RL methods.
Method Meta-World 50 Tasks
RLPD $7.62_{\pm 0.10}$
IBRL $17.08_{\pm 0.16}$
Cal-QL $24.60_{\pm 0.06}$
MTDIFF-P $27.06_{\pm 0.42}$
HarmoDT-F $57.37_{\pm 0.34}$
SODP (ours) $\textbf{59.26}_{\pm 0.18}$
@@ 5.4 Effectiveness of BC Regularization
Figure 5: Learning dynamics for different regularization. Performance declines without any regularization. Both KL and PL regularization confine the model to sub-optimal regions, while our BC regularization guides the model toward optimal actions.
@@
Figure 6: Visualization of trajectories using generated actions for different regularization. KL and PL regularization result in conservative policies, whereas our BC regularization preserves pre-trained knowledge while discovering new actions that lead to high rewards.
To demonstrate the effectiveness of our BC regularization, we conduct an ablation study by fine-tuning the same pre-trained model with our BC regularization and other regularization methods. We consider the following variants, each employing a different regularization strategy: (i) SODP w/o regularization. This variant fine-tunes the model directly using Eq. (3.2) without any regularization. (ii) SODP_kl. This variant uses a KL regularization term to constrain the divergence between the fine-tuned model and the pre-trained model. (iii) SODP_pl. This variant incorporates the original diffusion pre-training loss (PL) into the fine-tuning objective function. The details of these variants are provided in Appendix E, and additional ablation studies on different fine-tuning methods, such as directly fine-tuning with high-quality data, can be found in Appendix C.1.
Figure 5 demonstrates the effectiveness of our regularization in achieving a higher success rate. Directly fine-tuning the model without regularization leads to the worst performance, as the success rate declines due to the model degrading its pre-trained capabilities without constraints. However, adding KL and PL regularization is insufficient, as they cause oscillations near the pre-trained model. We hypothesize that the effectiveness of our BC regularization stems from two factors: (i) it enables the model to reuse acquired skills, preventing performance decline; (ii) it facilitates effective exploration of optimal regions by using optimal $\mu$ as the target policy. To validate this, we visualize trajectories generated by our planner’s actions using t-SNE (Van der Maaten & Hinton, 2008). As shown in Figure 6, after fine-tuning with KL and PL regularization, the trajectory distribution remains similar to the pre-training distribution, indicating that the model is reusing learned actions and lacks exploration into new regions. The exploration in PL is unstructured as it may lead to worse regions (e.g., the upper-left region in basketball). In contrast, our method demonstrates superior exploration capabilities to discover new, high-reward regions based on acquired knowledge (e.g., the lower-left region in basketball and the bottom region in plate-slide). Meanwhile, the model can derive valuable insights from pre-trained knowledge by exploiting discovered high-reward actions (e.g. the central region in plate-slide) while discarding low-reward actions.
@@ 5.5 Effectiveness of Pre-training
We investigate the impact of pre-training by comparing the performance of SODP with a variant that is fine-tuned from scratch using the same loss function defined in Eq. (15), denoted as SODP_scratch. Without pre-training, the policy is initially random, which leads to inaccurate estimation of the target policy $\mu$ in the BC regularization term. To mitigate this issue, we initialize the replay buffer with the same initial rollouts used during the fine-tuning of SODP. These rollouts are generated by the pre-trained model, enabling a more accurate approximation of the target policy $\mu$ during early fine-tuning.
We present the learning dynamics of the two methods in Figure 7. Directly fine-tuning the diffusion planner from scratch leads to worse performance. Without pre-training, the planner lacks an action prior to guide its behavior, resulting in stagnation as it struggles to explore high-reward regions. Moreover, the learning process becomes unstable, as the limited useful knowledge is easily overwritten by numerous ineffective trials. We further conduct ablation studies to examine the effect of varying the pre-training data, with results presented in Appendix C.2.
Benefits of Multi-Task Pre-Training.
To explore the benefits of pre-training on a multi-task dataset compared to a single-task dataset, we conduct experiments by fine-tuning the model on tasks excluded from the pre-training dataset. Specifically, we pre-train a model on the MT-10 dataset (SODP_mt10) and fine-tune it on three tasks not included in the pre-training data. We compare SODP_mt10 with a variant pre-trained exclusively on the basketball dataset (SODP_bas). As shown in Table 3, pre-training on multi-task data improves generalizability to unseen tasks as multi-task data provide a broader range of action distribution priors compared to single-task data.
Table 3: Average success rate on unseen tasks.
Unseen tasks SODP_mt10 SODP_bas
drawer-open $34.7_{\pm 0.06}$ $0.0_{\pm 0.0}$
plate-slide-side $55.3_{\pm 0.33}$ $0.0_{\pm 0.0}$
handle-pull-side $71.3_{\pm 0.13}$ $0.0_{\pm 0.0}$